answersLogoWhite

0

🎒

Statistics

Statistics deals with collecting, organizing, and interpreting numerical data. An important aspect of statistics is the analysis of population characteristics inferred from sampling.

36,755 Questions

How many pounds per square inch is high heel shoes?

You need to know two more pieces of information:

1. The weight of the person.

2. The area of the bottom of the heel contacting the floor.

If both feet are on the Floor, add the Area of BOTH Heels (in Sq Inches) together and divide the weight (in Pounds) by the Area (in Sq Inches). For Example:

Area = 1/2 Square Inch (1/4 + 1/4 Sq Inches)

Weight = 100 Pounds

Force on Floor = 100 lbs / (1/2 Sq In.)

= 200 lbs/Sq In.

How many numbers divisible by 3 or 4 are there from 1 to 20?

3=

3,6,9,12,15,18 (6 of them)

4=

4,8,12,16,20 (5 of them)

6+5=11

But only 1 number (12) divisibly by both.

What can you do a bar chart on?

you could do a bar chart on what types of crisps people like or on which days it rains.

How does the lack of a normal distribution influence the analytical options available to a researcher?

Most parametric and many non-parametric tests for testing the significance of hypotheses are based on the assumption of normality - or approximate normality.

Find the median of 4 6 9 6 7 6?

I've ranked you're data with the two centered values in parenthesis: 4 6 (6 6) 7 9, median = (6+6)/2 = 6

How you can find the brobability as a fraction?

I assume the probability is given a percent, or as odds, i.e. there's a 20% chance or 1 in 5 chance. You convert a percent by dividing by 100, 20% = 0.20, you convert odds by dividing the first number by the second, ie: 1/5 = 0.20.

How do you find the mode number?

It is the most commonnly occurring number in a set of numbers. I.E. if a set is 1,2,3,2,5,6,2 the mode is 2.

Why is it the heart so hard to draw?

It may be easy to draw if you use symmetry.........

What does it mean to say that the chances of getting a head when a coin is tossed are one-half?

It means that if the coin were tossed an infinite number of times,

half of the tosses would come up heads and half would be tails.

How did you get the answer of 45 if you are given the a question of The median of four numbers is 48 and the other 3 are 42 51 and 52?

The median of four numbers is the average of the second and third numbers when the four are in order. Suppose the median is m, and the missing number is x.

51 must be one of two middle numbers. So we have 48 = m = (x+51)/2 where x i sthe missing number.

96 = x+51 so x = 45

i would figure it out and show clear working and steps - i am not stupid at math and i took me about 5 minutes to figure out what you were doing.

we are trying to find x (45) to do thiswe have an equation using what we know that is relevant to finding the answer. we dont know 45 so we use x. and we know x + 51 ( the middle number we know) divided by 2 is 48

(x + 51) / 2 = 48

now we have to get x by itself on one side of the equation. so we should get rid of divide by 2 and you do that by multiply. and what you do to one side of the equation you must do to the other side to balance it out.

(x + 51) /2 x2 = 48 x2

now to simplify that

(x + 51) = 96

now we can just go 96 - 51 = x and x = 45

hope that was easier to understand!!

How many decimeters are in 83 meters?

1 m (meter) = 10 dm (decimeter)

1 m = 100 cm (centimeter)

1 m = 1000 mm (milimeter)